# On the Performance of Dual-Antenna Repeater Assisted Bi-Static MIMO ISAC

### v1

@@ I Introduction
Repeaters are active scatterers capable of receiving and retransmitting signals instantly with amplification. They can be deployed on a large scale to enhance coverage for multiple-input multiple-output (MIMO) wireless systems [Erik_WCL, Sara_Vieira_Erik_Mag, Jianan_TWC, Byoung_Magazine, Sergiy_WCL, Emil_repeater]. Reference [Sara_Vieira_Erik_Mag] demonstrated that repeaters can procure a spectral-efficiency (SE) similar to that of distributed MIMO, while obviating the front-haul signalling overhead. Furthermore, [RIS_Makki_IEEE_Acess] indicated that repeaters can outperform reconfigurable intelligent surface (RIS)-assisted systems, with the additional advantage of repeaters being band/frequency selective. The attractive implementation aspects of repeaters have led to their inclusion in the 5G NR standards since 3rd Generation Partnership Project (3GPP) Release $18$ [3gpp].
On parallel avenue, integrated sensing and communication (ISAC) offers efficient spectral and hardware utilization, and is actively being pushed forward as part of 3GPP Release $19$, a 5G-advanced/pre-6G initiative [ISAC_6G]. A detailed survey of ISAC is beyond the scope of this paper; readers can refer to [ISAC_survey, Zhang_Survey_2022, Michail_TWC] and references therein. In particular, the authors in [Erik_EUSIPCO] considered repeater-assisted mono-static ISAC  [^1] In a mono-static setting, the base station (BS) is required to be full-duplex, which in turn incurs additional signal processing and hardware overhead for self-interference cancellation., where the repeater was used to improve the coverage for downlink communication users. In contrast to [Erik_EUSIPCO], we investigate the feasibility of bi-static ISAC in the presence of a repeater [^2] Specifically, we consider a dual-antenna repeater (see [Sara_Vieira_Erik_Mag, Fig. $2$])., and focusing on the effects of the repeater on the performance of both target detection and downlink communication.
@@
However, the repeater also amplifies communication-to-sensing and sensing-to-communication interference. Specifically, the amplified sensing signal can incur increased interference for the communication users. Contrarily, for sensing, should the amplified downlink signals (directly via the repeater) overwhelm the reflected amplified echo of the target, detection performance at the sensing receiver can severely degrade. This paper investigates these tradeoffs and discusses the choice of precoder at the transmit BS that can mitigate this interference while retaining the ISAC performance improvement due to the repeater. Moreover, the developed analyses are agnostic of the channel models and applicable to diverse propagation conditions among the repeater, BS, and users.
@@
(b) Bi-static massive MIMO ISAC: sensing view-point. Clutters are not explicitly drawn to avoid congestion.
@@ II System Model
We consider a bi-static ISAC system where the transmit BS, with $N_{{\mathsf{t}}}$ antennas, beamforms a precoded sensing signal towards a target hot-spot area while simultaneously serving $K$ downlink single-antenna users. The system also contains a dual-antenna repeater. The receiver BS, with $N_{{\mathsf{t}}}$ antennas, performs target detection based on the reflected echoes from the target, also via the repeater. The system model from the communication point of view is illustrated in Fig. 1(a) and the sensing point of view in Fig. 1(b). The system entails several desired and interference channels, whose definitions are listed below.
@@
  • ${\mathbf{C}}{\in\mathbb{C}^{{N_{{\mathsf{r}}}\times N_{{\mathsf{t}}}}}}$: the composite clutter channel. Specifically, the clutter channel consisting of the LoS and NLoS paths caused by permanent obstacles can be measured beforehand and canceled [Behdad_Ozlem_TWC]. The unknown part corresponds to the NLoS paths caused by temporary obstacles, that are encased in $\mathbf{C}$}.
Apart from the clutter channel, we assume perfect channel state information (CSI)at the receiver for target detection as a necessary first step to understand the overall system performance with an intuitive and tractable formulation. Further, let $\nu\in\mathbb{C}$ be the repeater amplification factor. The target is characterized by the associated RCS, denoted by $\alpha$, which follows $\alpha{\sim\mathcal{CN}\left({0,\sigma_{{\mathtt{T}}}^{2}}\right)}$, $\sigma_{{\mathtt{T}}}^{2}$ being the RCS variance. The RCS is unknown at the receiver and needs to be estimated. Finally, we adopt a model wherein target detection is performed in a given hotspot, and a larger area can be covered by sending sensing signals to such hotspot areas serially [Behdad_Ozlem_TWC].
@@ III Signalling Scheme
${\mathbf{x}}[\tau]=$ $\sqrt{\rho}{\left\{\sum\limits_{n=1}^{K}\sqrt{\pi_{n}}{\mathbf{p}}_{n}s_{n}[\tau]+\sqrt{\pi_{{\mathtt{T}}}}{\mathbf{p}}_{{\mathtt{T}}}s_{{\mathtt{T}}}[\tau]\right\}},$
$\hskip 100.00015pt\tau=1,2,\ldots,\tau_{{\mathtt{L}}},$ (1)
@@
III-A Sensing with repeater
@@
The direct (amplified) interference due to the repeater, i.e., $\left\{\nu\boldsymbol{b}_{{\mathsf{r}}}\boldsymbol{b}_{{\mathsf{t}}}^{T}{\mathbf{x}}[\tau]\right\}$}, is incorporated as a part of the clutter. However, the composite channel can be pre-estimated (in LoS, it can even be fully characterized by angle of arrivals (AoAs) and angle of departure (AoD)) to cancel this interference, similar to inter-BS interference. [^4] This can also be cancelled by chosing the precoder for target ${\left({\mathbf{I}}_{N_{{\mathsf{t}}}}-\frac{1}{{\left\|\boldsymbol{b}_{{\mathsf{t}}}\right\|_{2}^{2}}}\boldsymbol{b}_{{\mathsf{t}}}^{*}\boldsymbol{b}_{{\mathsf{t}}}^{T}\right)}\boldsymbol{a}_{{\mathsf{t}}},$ leading to the direct link interference $\nu\boldsymbol{b}_{{\mathsf{r}}}\boldsymbol{b}_{{\mathsf{t}}}^{T}{\mathbf{p}}_{{\mathtt{T}}}=0$, where ${\mathbf{p}}_{{\mathtt{T}}}$ is the precoder intended for the target. The resulting $\dot{{\mathbf{C}}}\approx{\mathbf{C}}$ encases the effects of environmental scatterers. Later, with (4), we will develop the target detection framework.
III-B Downlink communication with repeater
@@
Next, based on (4) and (7), we present generalized likelihood ratio test (GLRT)-based target detection framework and downlink signal-to-interference-plus-noise ratio (SINR) analysis.
@@ IV RCS Estimation & Target Detection
where $\overline{{\mathbf{r}}}$, $\overline{{\mathbf{c}}}$, and $\overline{{\mathbf{w}}}$ are the concatenation of ${\mathbf{r}}[\tau]\alpha$, $\dot{{\mathbf{C}}}{\mathbf{x}}[\tau]$, and $\dot{{\mathbf{w}}}_{{\mathtt{B}}}[\tau]$, respectively. The null hypothesis $\mathcal{H}_{0}$ represents the case that there is no target in the sensing area, while the alternative hypothesis $\mathcal{H}_{1}$ represents the existence of the target. The joint RCS estimation, target-free channel (i.e. clutter) estimation, and the hypothesis testing problem can be written as
${\left\{\hat{\alpha},\hat{\dot{{\mathbf{c}}}},\hat{\mathcal{H}}\right\}}={\underset{{{\left\{{\alpha},{\dot{{\mathbf{c}}}},{\mathcal{H}}\right\}}}}{{\mathrm{arg\,max}}}}\quad p\left({\alpha,\dot{{\mathbf{c}}},{\mathcal{H}}\;|\;\overline{{\mathbf{y}}}_{\mathtt{BS}}}\right),$ (9)
@@
Now, only the parameter remains to be computed in the sensing noise covariance $\boldsymbol{\Sigma}_{{\mathsf{s}}}[\tau]$, which is evaluated (treating transmitted symbols as pilots) as
$\boldsymbol{\Sigma}_{{\mathsf{s}}}[\tau]$ $={\mathbb{E}\left[\dot{{\mathbf{w}}}_{{\mathtt{B}}}[\tau]\dot{{\mathbf{w}}}_{{\mathtt{B}}}^{H}[\tau]\right]}$
$={\mathbb{E}\left[\tilde{{\mathbf{G}}}_{{\mathtt{B}}}{\mathbf{x}}[\tau]{\mathbf{x}}^{H}[\tau]\tilde{{\mathbf{G}}}_{{\mathtt{B}}}^{H}\right]}+{\left\{{\big\lvert{\nu}\big\rvert}^{2}\sigma^{2}_{\mathtt{R}}\right\}}\boldsymbol{b}_{{\mathsf{r}}}\boldsymbol{b}_{{\mathsf{r}}}^{H}+\sigma^{2}_{\mathtt{BS}}{\mathbf{I}}_{N_{{\mathsf{r}}}}.$
@@ V Downlink SE Analysis
$y_{\mathtt{UE},n}=$ $\sqrt{\rho}\sqrt{\pi_{n}}\dot{{\mathbf{f}}}_{n}^{T}{\mathbf{p}}_{n}s_{n}+\sqrt{\rho}\sum\limits_{n^{\prime}=1}^{K}\sqrt{\pi_{n^{\prime}}}\dot{{\mathbf{f}}}_{n}^{T}{\mathbf{p}}_{n^{\prime}}s_{n^{\prime}}$
@@
${\mathtt{SINR}}_{n}=\frac{\rho\pi_{n}{\big\lvert{\dot{{\mathbf{f}}}_{n}^{T}{\mathbf{p}}_{n}}\big\rvert}^{2}}{\rho\sum\limits_{n^{\prime}=1}^{K}\pi_{n^{\prime}}{\big\lvert{\dot{{\mathbf{f}}}_{n}^{T}{\mathbf{p}}_{n^{\prime}}}\big\rvert}^{2}+\rho\pi_{{\mathtt{T}}}{\big\lvert{\dot{{\mathbf{f}}}_{n}^{T}{\mathbf{p}}_{{\mathtt{T}}}}\big\rvert}^{2}+{\mathbb{E}\left[{\big\lvert{\dot{w}_{\mathtt{UE}}}\big\rvert}^{2}\right]}},$
@@
Next, for the choice of the precoders, the transmitter can either consider the composite channel $\left\{\dot{{\mathbf{f}}}_{n}\right\}$} or be completely agnostic of the repeater and consider $\left\{{{\mathbf{f}}}_{n}\right\}$}. We consider regularized zero-forcing (ZF) precoding scheme for the communication user, in which case ${\mathbf{p}}_{n}$ becomes
${\mathbf{p}}_{n}=\epsilon_{n}{\left(\sum\limits_{n^{\prime}=1}^{K}\dot{{\mathbf{f}}}_{n^{\prime}}\dot{{\mathbf{f}}}_{n^{\prime}}^{H}+\zeta_{{\mathtt{ZF}}}{\mathbf{I}}_{N_{{\mathsf{t}}}}\right)}^{-1}\dot{{\mathbf{f}}}_{n},$
where $\epsilon_{n}$ is the normalizer to ensure ${\left\|{\mathbf{p}}_{n}\right\|_{2}^{2}}=1$ and $\zeta_{{\mathtt{ZF}}}>0$ is a regularizer. For ${\mathbf{p}}_{{\mathtt{T}}}$ we can consider one of the following:
  1. 1.
    Target centric: In this case, the transmitter beamforms towards the target, i.e., ${\mathbf{p}}_{{\mathtt{T}}}=\epsilon_{{\mathtt{T}}}\dot{{\mathbf{f}}}_{{\mathtt{T}}}$, with $\dot{{\mathbf{f}}}_{{\mathtt{T}}}=\boldsymbol{a}_{{\mathsf{t}}}$ and $\epsilon_{{\mathtt{T}}}$ is an appropriate normalizer.
  2. 2.
    Communication centric: We null the (potentially destructive) interference from the sensing signal to the users by projecting ${\mathbf{p}}_{{\mathtt{T}}}$ onto the nullspace of the subspace spanned by the users’ channel vectors. In this case, ${\mathbf{p}}_{{\mathtt{T}}}=\epsilon_{{\mathtt{T}}}{\left({\mathbf{I}}_{N_{{\mathsf{t}}}}-{\mathbf{U}}{\mathbf{U}}^{H}\right)}\dot{{\mathbf{f}}}_{{\mathtt{T}}}$, where $\mathbf{U}$} is a unitary matrix formed with the orthonormal columns that span the column space of $\left\{\dot{{\mathbf{f}}}_{1},\dot{{\mathbf{f}}}_{2},\ldots,\dot{{\mathbf{f}}}_{K}\right\}$}.

### v2

@@ I Introduction
Repeaters (also known as full-duplex relays) are active scatterers, capable of receiving and retransmitting signals instantly with amplification [Jianan_TWC]. They can be deployed on a large scale to enhance coverage for multiple-input multiple-output (MIMO) wireless systems [Erik_WCL, Sara_Vieira_Erik_Mag, Jianan_TWC, Byoung_Magazine, Sergiy_WCL, Emil_repeater]. Reference [Sara_Vieira_Erik_Mag] demonstrated that repeaters can procure a spectral-efficiency (SE) similar to that of distributed MIMO, while obviating the front-haul signalling overhead. Furthermore, [RIS_Makki_IEEE_Acess] indicated that repeaters can outperform reconfigurable intelligent surface (RIS)-assisted systems, with the additional advantage of repeaters being band/frequency selective. The attractive implementation aspects of repeaters have led to their inclusion in the 5G NR standards since 3rd Generation Partnership Project (3GPP) Release $18$ [3gpp].
On parallel avenue, integrated sensing and communication (ISAC) offers efficient spectral and hardware utilization, and is actively being pushed forward as part of 3GPP Release $19$, a 5G-advanced/pre-6G initiative [ISAC_6G]. A detailed survey of ISAC is beyond the scope of this paper; readers can refer to [ISAC_survey, Zhang_Survey_2022, Michail_TWC] and references therein, and specifically for bi-static ISAC[Wang_TWC, Xiao_TIT]. The authors in [Erik_EUSIPCO] considered repeater-assisted mono-static ISAC, [^1] In a mono-static setting, the base station (BS) is required to be full-duplex, which in turn incurs additional signal processing and hardware overhead for self-interference cancellation. where the repeater was used to improve the coverage for downlink communication users. In contrast to [Erik_EUSIPCO], we investigate the feasibility of bi-static ISAC in the presence of a repeater (specifically, a dual-antenna repeater as in [Sara_Vieira_Erik_Mag, Fig. $2$]), focusing on the effects of the repeater on the performance of both target detection and downlink communication.
@@
However, the repeater also amplifies communication-to-sensing and sensing-to-communication interference. Specifically, the amplified sensing signal can incur increased interference for the communication users. Contrarily, for sensing, should the amplified downlink signals (directly via the repeater) overwhelm the reflected amplified echo of the target, detection performance at the sensing receiver can severely degrade. To this end, this paper investigates these tradeoffs and discusses the choices of precoders at the transmit BS that can mitigate this interference while retaining the ISAC performance improvement due to the repeater. Moreover, the developed generalized likelihood ratio test (GLRT) framework is agnostic of the channel models and applicable to diverse propagation conditions among the repeater, BS, and users.
@@
(b) Bi-static massive MIMO ISAC: sensing view-point. Clutters are not explicitly drawn to avoid congestion. We neglect the multi-hop reflected channels for the sensing purpose because of the three/multi-fold cascaded path loss effect.
@@ II System Model
We consider a bi-static ISAC system where the transmit BS, with $N_{{\mathsf{t}}}$ antennas, beamforms a precoded sensing signal towards a target hot-spot area while simultaneously serving $K$ downlink single-antenna users. The system also contains a dual-antenna repeater. The receiver BS, with $N_{{\mathsf{t}}}$ antennas, performs target detection based on the reflected echoes from the target, also via the repeater. The system model from the communication point of view is illustrated in Fig. 1(a) and the sensing point of view in Fig. 1(b). The system entails several desired and interference channels, whose definitions are:
@@
  • ${\mathbf{C}}{\in\mathbb{C}^{{N_{{\mathsf{r}}}\times N_{{\mathsf{t}}}}}}$: the composite clutter channel. Specifically, the clutter channel consisting of the LoS and NLoS paths caused by permanent obstacles can be measured beforehand and canceled [Behdad_Ozlem_TWC]. The unknown part corresponds to the NLoS paths caused by temporary obstacles, which are encased in $\mathbf{C}$}.
Apart from the clutter channel, we assume perfect channel state information (CSI)at the receiver for target detection as a necessary first step to understand the overall system performance with an intuitive and tractable formulation. [^2] The composite channel via the target is $\alpha\boldsymbol{a}_{{\mathsf{r}}}\boldsymbol{a}_{{\mathsf{t}}}^{T}$. We assume only $\boldsymbol{a}_{{\mathsf{r}}}\boldsymbol{a}_{{\mathsf{t}}}^{T}$ to be known, as the BS beamforms to a specific sensing zone whose associated angle of arrival (AoA) and angle of departure (AoD) are pre-calibrated. However, $\alpha$ is unknown and will be estimated as a part of the GLRT later. Further, let $\nu\in\mathbb{C}$ be the repeater amplification factor. The target is characterized by the associated RCS, denoted by $\alpha$, which follows $\alpha{\sim\mathcal{CN}\left({0,\sigma_{{\mathtt{T}}}^{2}}\right)}$, $\sigma_{{\mathtt{T}}}^{2}$ being the RCS variance. The RCS is unknown at the receiver and needs to be estimated. Finally, we adopt a model wherein target detection is performed in a given hotspot, and a larger area can be covered by sending sensing signals to such hotspot areas serially [Behdad_Ozlem_TWC].
@@ III Signalling Scheme
${\mathbf{x}}[\tau]=$ $\sqrt{\rho}{\left\{\sum\limits_{n=1}^{K}\sqrt{\pi_{n}}{\mathbf{p}}_{n}s_{n}[\tau]+\sqrt{\pi_{{\mathtt{T}}}}{\mathbf{p}}_{{\mathtt{T}}}s_{{\mathtt{T}}}[\tau]\right\}},~\tau\in{\left[\tau_{{\mathtt{L}}}\right]},$ (1)
@@
III-1 Sensing with repeater
@@
The direct (amplified) interference due to the repeater, i.e., $\left\{\nu\boldsymbol{b}_{{\mathsf{r}}}\boldsymbol{b}_{{\mathsf{t}}}^{T}{\mathbf{x}}[\tau]\right\}$}, is incorporated as a part of the clutter. However, the composite channel can be pre-estimated (in LoS, it can even be fully characterized by AoAs and AoDs) to cancel this interference, similar to inter-BS interference. [^4] This can also be cancelled by choosing the precoder for target ${\left({\mathbf{I}}_{N_{{\mathsf{t}}}}-\frac{1}{{\left\|\boldsymbol{b}_{{\mathsf{t}}}\right\|_{2}^{2}}}\boldsymbol{b}_{{\mathsf{t}}}^{*}\boldsymbol{b}_{{\mathsf{t}}}^{T}\right)}\boldsymbol{a}_{{\mathsf{t}}},$ leading to the direct link interference $\nu\boldsymbol{b}_{{\mathsf{r}}}\boldsymbol{b}_{{\mathsf{t}}}^{T}{\mathbf{p}}_{{\mathtt{T}}}=0$, where ${\mathbf{p}}_{{\mathtt{T}}}$ is the precoder intended for the target. Although it is not necessary for the subsequent analyses. The resulting $\dot{{\mathbf{C}}}\approx{\mathbf{C}}$ encases the effects of environmental scatterers. Later, with (4), we will develop the target detection framework.
III-2 Downlink communication with repeater
@@
Next, based on (4) and (7), we present GLRT-based target detection framework and downlink signal-to-interference-plus-noise ratio (SINR) analysis.
@@ IV RCS Estimation & Target Detection
The complexity of the GLRT is dominated by a matrix inversion of dimension $(N_{{\mathsf{t}}}N_{{\mathsf{r}}}+1)\times(N_{{\mathsf{t}}}N_{{\mathsf{r}}}+1)$, and hence is $\mathcal{O}\left((N_{{\mathsf{t}}}N_{{\mathsf{r}}}+1)^{3}\right)$. Next, only the parameter remains to be computed in the sensing noise covariance $\boldsymbol{\Sigma}_{{\mathsf{s}}}[\tau]={\mathbb{E}\left[\dot{{\mathbf{w}}}_{{\mathtt{B}}}[\tau]\dot{{\mathbf{w}}}_{{\mathtt{B}}}^{H}[\tau]\right]}$, which is evaluated (treating transmitted symbols as pilots) as
$\boldsymbol{\Sigma}_{{\mathsf{s}}}[\tau]={\mathbb{E}\left[\tilde{{\mathbf{G}}}_{{\mathtt{B}}}{\mathbf{x}}[\tau]{\mathbf{x}}^{H}[\tau]\tilde{{\mathbf{G}}}_{{\mathtt{B}}}^{H}\right]}+{\left\{{\big\lvert{\nu}\big\rvert}^{2}\sigma^{2}_{\mathtt{R}}\right\}}\boldsymbol{b}_{{\mathsf{r}}}\boldsymbol{b}_{{\mathsf{r}}}^{H}+\sigma^{2}_{\mathtt{BS}}{\mathbf{I}}_{N_{{\mathsf{r}}}}.$
@@
Remark 2.
We note that a noisy repeater can severely degrade the sensing performance, which is captured by ${\big\lvert{\nu}\big\rvert}^{2}\sigma^{2}_{\mathtt{R}}$ in the sensing noise covariance. The effect of the amplified downlink signals (directly via the repeater) is implicitly captured in the clutter covariance matrix in (11c).
@@ V Downlink SE Analysis
$y_{\mathtt{UE},n}=$ $\sqrt{\rho}\sqrt{\pi_{n}}\dot{{\mathbf{f}}}_{n}^{T}{\mathbf{p}}_{n}s_{n}+\sqrt{\rho}\sum\limits_{n^{\prime}=1,n^{\prime}\neq n}^{K}\sqrt{\pi_{n^{\prime}}}\dot{{\mathbf{f}}}_{n}^{T}{\mathbf{p}}_{n^{\prime}}s_{n^{\prime}}$
@@
${\mathtt{SINR}}_{n}=\frac{\rho\pi_{n}{\big\lvert{\dot{{\mathbf{f}}}_{n}^{T}{\mathbf{p}}_{n}}\big\rvert}^{2}}{\rho\sum\limits_{n^{\prime}=1,n^{\prime}\neq n}^{K}\pi_{n^{\prime}}{\big\lvert{\dot{{\mathbf{f}}}_{n}^{T}{\mathbf{p}}_{n^{\prime}}}\big\rvert}^{2}+\rho\pi_{{\mathtt{T}}}{\big\lvert{\dot{{\mathbf{f}}}_{n}^{T}{\mathbf{p}}_{{\mathtt{T}}}}\big\rvert}^{2}+{\mathbb{E}\left[{\big\lvert{\dot{w}_{\mathtt{UE}}}\big\rvert}^{2}\right]}},$
@@
Next, for the choice of the precoders, the transmitter can either consider the composite channel $\left\{\dot{{\mathbf{f}}}_{n}\right\}$} or be completely agnostic of the repeater and consider $\left\{{{\mathbf{f}}}_{n}\right\}$}. We consider regularized zero-forcing (ZF) precoding scheme for the communication user, in which case ${\mathbf{p}}_{n}$ becomes ${\mathbf{p}}_{n}=\epsilon_{n}{\left(\sum\limits_{n^{\prime}=1}^{K}\dot{{\mathbf{f}}}_{n^{\prime}}\dot{{\mathbf{f}}}_{n^{\prime}}^{H}+\zeta_{{\mathtt{ZF}}}{\mathbf{I}}_{N_{{\mathsf{t}}}}\right)}^{-1}\dot{{\mathbf{f}}}_{n},$ where $\epsilon_{n}$ is the normalizer to ensure ${\left\|{\mathbf{p}}_{n}\right\|_{2}^{2}}=1$ and $\zeta_{{\mathtt{ZF}}}>0$ is a regularizer. For ${\mathbf{p}}_{{\mathtt{T}}}$ we can consider one of the following: Target centric: In this case, the transmitter beamforms towards the target, i.e., ${\mathbf{p}}_{{\mathtt{T}}}=\epsilon_{{\mathtt{T}}}\dot{{\mathbf{f}}}_{{\mathtt{T}}}$, with $\dot{{\mathbf{f}}}_{{\mathtt{T}}}=\boldsymbol{a}_{{\mathsf{t}}}$ and $\epsilon_{{\mathtt{T}}}$ is an appropriate normalizer. Communication centric: We null the (potentially destructive) interference from the sensing signal to the users by projecting ${\mathbf{p}}_{{\mathtt{T}}}$ onto the nullspace of the subspace spanned by the users’ channel vectors. In this case, ${\mathbf{p}}_{{\mathtt{T}}}=\epsilon_{{\mathtt{T}}}{\left({\mathbf{I}}_{N_{{\mathsf{t}}}}-{\mathbf{U}}{\mathbf{U}}^{H}\right)}\dot{{\mathbf{f}}}_{{\mathtt{T}}}$, where $\mathbf{U}$} is a unitary matrix formed with the orthonormal columns that span the column space of $\left\{\dot{{\mathbf{f}}}_{1},\dot{{\mathbf{f}}}_{2},\ldots,\dot{{\mathbf{f}}}_{K}\right\}$}.